The length of a rectangular sign is 3 feet longer than the width. If the sign has an area of 40 square feet for advertising, find its length and width.
Can someone show me how to do this?

Answer by checkley71(8403)

area=length times width or 40=x(x+3) or 40=x~2+3x or x~2+3x-40=0 or
(x+8)(x-5)=0 or x=-8 & x=5 so the sides are 5 & 5+3=8